# Actor-Critic Physics-informed Neural Lyapunov Control

### v1

@@ III-E Actuation Constraints
$\pi_{\gamma}(x)\!=\!V\cdot\mathrm{softmax}(u_{\gamma}(x))\!=\!\sum_{i=1}^{M}V_{i}\cdot\mathrm{softmax}(u_{\gamma}(x))_{i}$ (13)
where the vector-valued softmax function simply generates the coefficients of the convex combination of columns of $V$ and $u_{\gamma}(\cdot)$ is a neural network. In contrast to (12), this parameterization eliminates the need to compute any numerical optimization sub-routine.
@@ III-F Verification
By minimizing the empirical loss over the sampled states, the learned controller is not guaranteed to be stabilizing, thus we need to formally verify the learned certificate function.
Let $c$ be a real number between 00 and $1$, we attempt to verify the following three conditions
$\begin{cases}W_{\theta}(x)>W_{\theta}(0)~{}\forall x\in R_{2},W_{\theta}(x)<c,\|x\|\geq\epsilon\\ \dot{W}_{\theta}(x)<0~{}\forall x\in R_{2},W_{\theta}(x)<c,\|x\|\geq\epsilon\\ W_{\theta}(x)>c~{}\forall x\in\partial R_{2}\end{cases}$ (14)
Because of the complexity of neural network functions, a sublevel set may consist of multiple disjoint sets, we only consider the sublevel set in the region of interest. The first and second conditions verify the Lyapunov condition on the set $\mathcal{D}=R_{2}\cap\{x:W_{\theta}(x)<c\}$. We ignore a smaller neighbor around the origin to avoid numerical errors. However, the first two conditions do not imply that $\mathcal{D}$ is the domain of attraction. We also need to verify the third condition, which implies that the set $\mathcal{D}$ is strictly contained in $R_{2}$ so that a trajectory does not leave $R_{2}$ with negative $\dot{W}_{\theta}$ along the trajectory. If these two conditions are satisfied, then the function $W_{\theta}(x)-W_{\theta}(0)$ is a Lyapunov function in the region $R_{2}\cap\{x:W_{\theta}(x)<c\}$.
@@ IV Numerical Experiments
We demonstrate the effectiveness of our method on various nonlinear control problems. We first run our method on Double Integrator and Van der Pol and compare the verified DoA obtained by our method with that of the LQR controller. We then test our method on Inverted Pendulum and Bicycle Tracking and compare it to both LQR and Neural Lyapunov Control [10]. Across all examples, we use the same hyperparameter $M=8,K=64,\lambda_{0}=5,\lambda_{c}=0.5$, and $\lambda_{b}=5$ for the loss function. Other hyperparameters can be found in Table I. In all experiments, we assumed that the actuation constraint is $u\in\mathcal{U}=\{u|~{}\|u\|_{\infty}\leq 1\}$ and the learned controller satisfies this constraint by using $\mathrm{tanh}$ as the activation function in the output layer.
To find the DoA of LQR under actuation constraints, we first solve the Riccati equation of the linearized system with $Q=I,R=I$ to obtain the solution $P$ and control matrix $K$, and then we verify the following conditions
$x^{\top}P\dot{x}<0\text{ and }\|Kx\|_{\infty}\leq 1~{}\forall x^{\top}Px<c_{lqr},\|x\|\geq\epsilon$ (15)
by an SMT solver, where $\dot{x}$ is given by the nonlinear closed-loop dynamics. To find the largest DoA, we perform bisection to find the largest $c_{lqr}$ such that both conditions are satisfied.
We used the SMT solver dReal [28] to verify the conditions (14) for the neural network controller and (15) for the LQR controller.
@@ IV-A Double Integrator
We first consider the double integrator dynamics,
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ u\end{bmatrix}$
The learned Lyapunov function is shown in Figure 1(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 1(b).
@@ IV-B Van der Pol
Consider the Van der Pol system
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ x_{1}-\mu(1-x_{1}^{2})x_{2}+u\end{bmatrix}$
The learned Lyapunov function is shown in Figure 2(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 2(b).
@@ IV-C Inverted Pendulum
The inverted pendulum system has two states, the angular position $\theta$, angular velocity $\dot{\theta}$ and control input $u$. The dynamical system of inverted pendulum is
$\begin{bmatrix}\dot{\theta}\\ \ddot{\theta}\end{bmatrix}=\begin{bmatrix}\dot{\theta}\\ \frac{g}{l}\sin(\theta)-\frac{b\dot{\theta}}{ml^{2}}+\frac{u}{ml^{2}}\end{bmatrix}$
The learned Lyapunov function is shown in Figure 3(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 3(b).
@@ IV-D Bicycle Tracking
The bicycle tracking system has two states, the distance $d_{e}$, angle error $\theta_{e}$, and control input $u$. The dynamical system for the tracking problem is
$\begin{bmatrix}\dot{d}_{e}\\ \dot{\theta}_{e}\end{bmatrix}=\begin{bmatrix}v\sin(\theta_{e})\\ v\frac{\tan(u)}{l}-\frac{\cos(\theta_{e})}{1-d_{e}}\end{bmatrix}$
The learned Lyapunov function is shown in Figure 4(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 4(b).
@@
All the experiments show that the verified DoA obtained by our method is 2$\sim$4 times larger than all other methods.
@@
TABLE I: Hyperparameters for each experiment
@@ V CONCLUSIONS
In this paper, we developed an actor-critic framework to jointly train a neural network control policy and the corresponding stability certificate, with the explicit goal of maximizing the induced region of attraction. To this end, we leveraged Zubov’s partial differential equation to inform the loss function of the shape of the boundary of the true domain of attraction. Our numerical experiments on several nonlinear benchmark examples corroborate the superiority of the proposed method over competitive approaches in enlarging the domain of attraction. For future work, we will incorporate robustness to model uncertainty in our loss function. We will also investigate the extension of the proposed method to discrete-time systems [16].

### v2

@@ III-E Actuation Constraints
$\pi_{\gamma}(x)\!=\!V\cdot\mathrm{softmax}(u_{\gamma}(x))\!=\!\sum_{i=1}^{M}\!V_{i}\mathrm{softmax}(u_{\gamma}(x))_{i}$ (13)
where the vector-valued softmax function simply generates the coefficients of the convex combination of columns of $V$. In contrast to (12), this parameterization eliminates the need to compute any numerical optimization sub-routine.
@@ III-F Verification
To formally verify the learned certificate, let $c\in(0,1)$. Our goal is to verify the following three conditions
$\begin{cases}W_{\theta}(x)>W_{\theta}(0)~{}\forall x\in\{x\in R_{2}\mid W_{\theta}(x)<c,\ \|x\|\geq\epsilon\}\\ \dot{W}_{\theta}(x)<0~{}\forall x\in\{x\in R_{2}\mid W_{\theta}(x)<c,\ \|x\|\geq\epsilon\}\\ W_{\theta}(x)>c~{}\forall x\in\partial R_{2}\end{cases}$ (14)
Since $W_{\theta}$ can have disjoint sub-level sets, we only consider sub-level sets contained in $R_{2}$.
The first and second conditions verify the Lyapunov conditions on the set $\mathcal{D}=R_{2}\cap\{x:W_{\theta}(x)<c\}$. We ignore a smaller neighbor around the origin to avoid numerical errors. However, the first two conditions do not imply that $\mathcal{D}$ is the domain of attraction. We also need to verify the third condition, which implies that the set $\mathcal{D}$ is strictly contained in $R_{2}$ so that a trajectory does not leave $R_{2}$ with negative $\dot{W}_{\theta}$ along the trajectory. If these conditions are satisfied, then the function $W_{\theta}(x)-W_{\theta}(0)$ is a Lyapunov function in the region $R_{2}\cap\{x:W_{\theta}(x)<c\}$.
@@ IV Numerical Experiments
We demonstrate the effectiveness of our method on various nonlinear control problems. We first run our method on Double Integrator and Van der Pol and compare the verified DoA obtained by our method with that of the LQR controller. We then test our method on Inverted Pendulum and Bicycle Tracking and compare it to both LQR and Neural Lyapunov Control [13] [^1] The code repository of [13] does not support the first two examples.. Across all examples, we use the same hyperparameter $M=8,K=64,\lambda_{0}=5,\lambda_{c}=0.5$, and $\lambda_{b}=5$ for the loss function. Other hyperparameters and verified sublevel set can be found in Table I. The effect of different choices of hyperparameters can be found in the preprint version [31]. In all experiments, we assumed that the actuation constraint is $u\in\mathcal{U}=\{u|~{}\|u\|_{\infty}\leq 1\}$ and the learned controller satisfies this constraint by using $\mathrm{tanh}$ as the activation function in the output layer.
To find the DoA of LQR under actuation constraints, we first solve the Riccati equation of the linearized system with $Q=I,R=I$ to obtain the solution $P$ and control matrix $K$, and then we verify the following conditions,
$\begin{cases}x^{\top}P\dot{x}<0~{}\forall x\text{ s.t }x^{\top}Px<c_{lqr}\text{ and }\|x\|\geq\epsilon\\ \|Kx\|_{\infty}\leq 1~{}\forall x\text{ s.t }x^{\top}Px<c_{lqr}\text{ and }\|x\|\geq\epsilon,\end{cases}$ (15)
where $\dot{x}$ is given by the nonlinear closed-loop dynamics.
We used the SMT solver dReal [32] to verify the conditions (14) for the neural network controller and (15) for the LQR controller. To find the largest DoA for both LQR and our method, we perform bisection to find the largest $c_{lqr}$ ($c$) such that conditions (14), (15) are satisfied.
@@ IV-A Double Integrator
We first consider the double integrator dynamics $\dot{x}_{1}=x_{2},\ \dot{x}_{2}=u$. The learned Lyapunov function is shown in Figure 2(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 2(b).
@@ IV-B Van der Pol
Consider the Van der Pol system $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=x_{1}-\mu(1-x_{1}^{2})x_{2}+u$. The learned Lyapunov function is shown in Figure 3(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 3(b).
@@ IV-C Inverted Pendulum
The inverted pendulum system has two states, the angular position $\theta$, angular velocity $\omega$, and control input $u$. The dynamical system of inverted pendulum is $\dot{\theta}=\omega$ and $\dot{\omega}=\frac{g}{l}\sin(\theta)-\frac{b\omega}{ml^{2}}+\frac{u}{ml^{2}}$. The learned Lyapunov function is shown in Figure 4(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 4(b).
@@ IV-D Bicycle Tracking
The bicycle tracking system has two states, the distance $d_{e}$, angle error $\theta_{e}$, and control input $u$. The dynamical system for the tracking problem is $\dot{d}_{e}=v\sin(\theta_{e})$ and $\dot{\theta}_{e}=v\frac{\tan(u)}{l}-\frac{\cos(\theta_{e})}{1-d_{e}}$. The learned Lyapunov function is shown in Figure 5(a). The verified DoA and the vector field induced by the learned controller are shown in Figure 5(b).
@@
In summary, all the experiments show that our verified DoA is 2$\sim$4 times larger than other methods.
@@
TABLE I: Experiment Details
@@ V CONCLUSIONS
In this paper, we developed an actor-critic framework to jointly train a neural network control policy and the corresponding stability certificate, with the explicit goal of maximizing the induced region of attraction by leveraging Zubov’s partial differential equation to inform the loss function of the shape of the boundary of the true domain of attraction. Our numerical experiments on several nonlinear benchmark examples corroborate the superiority of the proposed method over competitive approaches in enlarging the domain of attraction. For future work, we will incorporate robustness to model uncertainty in our loss function. We will also investigate the extension of the proposed method to discrete-time systems [20].